OpenStudy (anonymous):

when n=14,206 & 957 show that sigma(n)=sigma(n+1)

OpenStudy (anonymous):

\[\sigma(14)=\sigma(2*7)=24=\sigma(3*5)=\sigma(15)\]

OpenStudy (anonymous):

is it 2*7 because that is the prime factorization? then i assume you would \[2^2-1/2-1*7^2-1/7-1 \] which is 8? not 24

ganeshie8 (ganeshie8):

\(\sigma(14) = \sigma(2) \sigma(7) = (1+2)(1+7) = 24\)

OpenStudy (anonymous):

oh okay so you take sigma of each individual prime factor

OpenStudy (anonymous):

and how does that equal to (3*5)?

ganeshie8 (ganeshie8):

yes cuz \(\sigma\) is a multiplicative function : \(\sigma(ab) = \sigma(a) \sigma(b)\)

OpenStudy (anonymous):

im sorry one more question how is that equal to 3*5?

ganeshie8 (ganeshie8):

\(\sigma(15) = \sigma(3) \sigma(5) = (1+3)(1+5) = 24\)

OpenStudy (anonymous):

okay got it thanks!